OpenStudy (bananas1234):

Simplify. a) (3xy3)(2x3y) b) 2y(2y + 3y2) – y2(3 – y)

OpenStudy (bee_see):

For a) separate them into different parts so (3)(2) (x)(x^3) (y^3 )(y)

OpenStudy (bee_see):

If you have something like a^n times a^m then a^n+m

OpenStudy (bananas1234):

like this?.....3*2 =6 and x*x^3 = x^(1+3) = x^4 and y^3 *y = y^(3+1) = y^4

OpenStudy (bee_see):

Yes! Now just put them together.

OpenStudy (bee_see):

Like this: 6x^4y^4

OpenStudy (bananas1234):

3*2 = 6x*x^3 = x^(1+3) = x^4 y^3 *y = y^(3+1) = y^4

OpenStudy (bee_see):

Put all the parts together. The answer will look like this: 6x^4y^4.

OpenStudy (bananas1234):

so for the first part do i multiply 3 times 2 and so on?

OpenStudy (bee_see):

Yes, you did it fine the first time. You just have to take all 3 parts and put them together. Didn't you get, 6, x^4, and y^4? Just put them all together. 6x^4y^4. That's the answer.

OpenStudy (bananas1234):

oh i think i get that now :)

OpenStudy (bee_see):

now for b, you have to distribute.

OpenStudy (bananas1234):

ok

OpenStudy (bee_see):